54-6 divided by 2 +2=? I just want to make sure I understand the answer, you start by dividing, then subtract 54 - the answer by dividing (3) then add the 6 which gives you 57
Answer by checkley77(12569)

YOU NEED TO USE ( ) IF YOU WANT US TO PROPERLY INTERPRET YOUR PROBLEM BECAUSE THERE ARE TOO MANY OPTIONS AS SHOWN BELOW.
(54-6)/(2+2)=48/4=12
54-(6/2)+2=54-3+2=53
54-6/(2+2)=54-6/4=54-1.5=52.5
(54-6)/2+2=48/2+2=24+2=26